27. A researcher claims that the number of motor vehicle crash deaths is consistent across age groups. The table shows the results of a random sample of motor vehicle crash deaths by age group. Are the numbers of crash deaths significantly different by age group, using a 0.05 level of significance? (15 points – Show your work!)

Age Group:

16-24

25-34

35-44

45-54

55-64

65-74

75 and older

Total

# of Crash Deaths

135

131

97

109

100

65

63

700


What are your hypotheses?


H0:

HA:

Fill in the following table to calculate your test statistic:

Age Group:

16-24

25-34

35-44

45-54

55-64

65-74

75 and older

Total

# of Crash Deaths

135

131

97

109

100

65

63

700

# Expected

(O – E)

(O – E)2 / E

df =______

Critical value: _________________


Conclusion: We ________ (reject / fail to reject) the Null Hypothesis

Interpretation:
Answer #1

Ho:  numbers of crash deaths are not significantly different by age group

H1:  numbers of crash deaths are significantly different by age group

Chi square test for Goodness of fit  
  
expected frequncy,E = expected proportions*total frequency = 1/7*700  
total frequency=   700

observed frequencey, O expected proportion expected frequency,E (O-E) (O-E)²/E
135 0.143 100.00 35.00 12.250
131 0.143 100.00 31.00 9.610
97 0.143 100.00 -3.00 0.090
109 0.143 100.00 9.00 0.810
100 0.143 100.00 0.00 0.000
65 0.143 100.00 -35.00 12.250
63 0.143 100 -37.00 13.690

chi square test statistic,X² = Σ(O-E)²/E =   48.700

level of significance, α=   0.05              
Degree of freedom=k-1=   7   -   1   =   6
Critical value =    12.592 [ Excel function: =chisq.inv.rt(α,df) ]      

since, test stat > critical value, reject Ho

We (reject ) the Null Hypothesis

Interpretation: we have enough evidence to conclude that numbers of crash deaths significantly different by age group